Department of Mathematics University of Utah

COURSE OUTLINE FOR MATHEMATICS 5710 FOR FALL 2001

Time and Place: 10.45-11.35 MWF, JWB 208 Instructor: Graeme Milton Office: 310 JWB Phone: 581-6495 Email: milton@math.utah.edu Office Hours: I’ll try to arrange office hours to suit the majority of people. However

if this scheduled time is unsuitable for you, please call or email me to make a separate appointment. You could also try dropping by my office, but please not right before class.

Text: Strang, Introduction to Applied Mathematics, Wellesley-Cambridge Press (1986) \underline{Content of the Course}: The course for this semester is based on Chapters 1, 2, 3 and 8 of

the text, but will also include some supplementary material not in the text. The topics to be covered during the semester include: Symmetric linear systems, positive definite matrices, eigenvalue problems, equilibrium equations for discrete and continuous systems, the Kalman filter, boundary value problems in ODEs and PDEs and boundary integrals, optimization, linear programming, and game theory.

\underline{Prerequisites}: The recommended prerequisite for Mathematics 5710 is Mathematics 2250. The courses 3150 and 3160 are NO LONGER prerequisites.

\underline{Reading and Problems}: The classes are \underline{not} a substitute for studying the text and working the assigned problems.

\underline{Assignments} If you wish you may collaborate on the assignments. You should however each hand in a separate assignment for grading. The assignments will be given out about once a week. Only a subset of the assigned problems will be graded, but you should try to do all of the assigned problems (similar questions may be on the exams.)

\underline{Grading Policy}: Grades will be based on the assignments (40% of the grade), on a midterm (25% of the grade) and on a final (35% of the grade).

\underline{Exams}: The mid-term will be held on Friday October 12th, during the usual scheduled class time. The final exam will be held on Wednesday December 12th, 10:30 a.m. - 12:30 p.m. These are closed book examinations. You may use hand held personal calculators. The use of laptop computers will not be permitted.

\underline{Late homework, missed exams} : You can hand in assignments late, without penalty, provided I have not started grading them. However thats a gamble, as you will get zero credit if I have started grading them and I’ll try to grade promptly. If you are sick, let me know not later than the day assignments are due and I’ll delay the grading. You can retake a missed exam only for reasons of medical illness substantiated by a letter from your doctor.

Assignment 1, Due Friday August 31st

Section1.2 problems 1.2.1, 1.2.10∗, 1.2.12∗, 1.2.13
Section1.3 problems 1.3.1, 1.3.11∗, 1.3.12∗, 1.3.14

The problems marked * will be graded and each of these is worth 2 points.